1. Saponification is a process wherein soap is added to a certain type of oil to produce grease. One

such grease is said to have 75.7 % by volume oil and 24.3% by volume soap, wherein the oil and

soap have densities 760 kgm/m3 and 6,250 kgm/m3, respectively. This kind of grease is sold by

packs shaped like a sphere with a diameter of 20 cm for convenience. Calculate: a) the mass of

oil per pack (kgm), b) the mass of soap per pack (kgm),c) the weight of each pack (kgf), d) the

specific weight of grease (N/m3), e) will the grease float if it is placed in water whose density is 1000 kgm/m3? Explain


1
Expert's answer
2022-02-21T16:09:37-0500

a)

"m_o=\eta_o\rho_oV=2.41~kg,"

b)

"m_s=\eta_s\rho_sV=6.36~kg,"

c)

"m=m_o+m_s=8.77~kg,"

d)

"\dot P=\frac PV=\frac{mg}V=20.5~\frac{kN}{m^3},"

e)

"\rho=\frac mV=2100~\frac{kg}{m^3}," the grease will not float.